# Sublinear Regret for a Class of Continuous-Time Linear–Quadratic Reinforcement Learning Problems

### v4

@@ 1 Introduction
Recently, there has been some progress on regret analysis for continuous-time stochastic LQ RL in (Basei et al., , 2022; Szpruch et al., , 2024) that achieves sublinear regrets of their respective algorithms. Both papers assume that the diffusion coefficients are constant (independent of state and control), which is vital for their approaches to work. Again, in essence, these works are model-based because they apply either least-square or Bayesian methods to estimate model parameters and use the corresponding estimation errors to deduce the regret bounds. In particular, there is an intrinsic caveat of this approach applied to LQ problems: when estimating the model parameters, an unidentifiability issue may arise (from optimal control policies being linear feedbacks of the state) and a special exploratory stage has to be carefully designed to avoid it; see e.g., Szpruch et al., (2021); Basei et al., (2022); Szpruch et al., (2024); Xu et al., (2024). In addition, to get one estimate of the model parameters, the required number of episodes and that of timesteps within each episode increase exponentially over the learning process, adding substantial computational and memory costs. [^1] For example, given a total $N$ episodes of the learning budget, the algorithm in Basei et al., (2022) suggests to break it into several epoches, with the $k$-th epoch containing $2^{k}$ episodes and $2^{k/2}$ timesteps and to update a policy after having applied it for an epoch of episodes. As a result, although the total regret is proved to be $O(\log N)$, the storage and computational complexity increase exponentially in the number of epoches (which is the number of times of updating the policy parameters).
@@ 2.1 Classical Stochastic LQ Control
We remark that if $x$ is of a higher dimension and/or there is a control running reward, then the optimal policy will depend explicitly on the solution of the differential Riccati equation and hence become time-dependent (Yong and Zhou, , 1999, Chapter 6), for which our current method fails.  [^2] For example, when $x$ is higher dimensional, the optimal control is $u_{CL}(t,x)=-(\sum_{j=1}^{m}D_{j}^{\top}\frac{\partial V_{CL}(t,x)}{\partial xx}D_{j})^{-1}\left(B^{\top}\frac{\partial V_{CL}(t,x)}{\partial x}+\sum_{j=1}^{m}D_{j}^{\top}\frac{\partial V_{CL}(t,x)}{\partial xx}C_{j}x\right)$, which becomes time-dependent.
@@ 2.2 RL Theory for LQ
where $\bar{\mu}=-(\sum_{j=1}^{m}D_{j}D_{j}^{\top})^{-1}(B+\sum_{j=1}^{m}C_{j}D_{j})$ and $\Sigma=\frac{\gamma}{k_{1}(t)}(\sum_{j=1}^{m}D_{j}D_{j}^{\top})^{-1}$.333 Detailed derivations of these results follow analogously from those in Wang et al., (2020) for the infinite horizon case. Note that these theoretical results cannot be used to compute the solution of the exploratory problem because the model parameters are unknown, yet they reveal the structure of the solution inherent to LQ RL (i.e. the optimal value function is quadratic in $x$ and optimal stochastic policy is Gaussian) that can be utilized to significantly reduce the complexity of function parameterization/approximation in learning.
@@ 3 A Continuous-Time RL Algorithm
This section presents the steps of designing a continuous-time RL algorithm for solving our LQ problem, including function parameterization, policy evaluation and policy gradient. [^4] As will be explained below, policy evaluation is actually not necessary for the specific LQ problem considered in this paper. However, we still include it in the discussion and in the algorithm for future extension to general problems where policy evaluation is generally needed and indeed a crucial step. We will introduce various techniques such as exploration scheduling and projection for deriving the convergence rate of the policy parameter and the regret bound. We will also describe time discretization for final implementation.
@@ 3.1 Function Parameterization
Note that (2.12) suggests that the optimal feedback policy is time-dependent, whose variance depends explicitly on $t$. In our parameterization, the time-dependent variance of the Gaussian policies is replaced by a decaying schedule, called an exploration schedule, of $\phi_{2}$ as a function of the number of iterations, to be presented shortly. [^5] An alternative approach is to set a temperature decaying schedule, or equivalently take $\phi_{2}$ as an endogenous, learnable parameter based on the expression (2.12). For example, Szpruch et al., (2024) proposes two different temperature scheduling sequences for their model-based algorithms respectively, which determine the variances of the respective policies in the learning processes that eventually impact the regrets of the algorithms. In this paper, we choose to take a shortcut by taking the entire policy variance as a suitable exogenous tuning sequence (as shown in Theorem 4.1), leaving the temperature $\gamma$ as a constant for simplicity.
@@ 3.5 Discretization
Our approach for continuous-time RL is characterized by carrying out the entire analysis in the continuous-time setting and discretizing time only at the final implementation stage. There are two reasons for discretization. First, note that in the updating rules (3.8) to (3.10), we implicitly assume a continuous sampling of control $u_{n}(t)|x_{n}(t)\sim\mathcal{N}(\phi_{1,n}x_{n}(t),\phi_{2,n})$, which is not practical for actual inplementation. Second, The iterations in (3.8) and (3.9) involve integrals that can be computed only by approximated discretized summations as well as the $\mathrm{d}J$ term that can be approximated by the temporal difference between two consecutive time steps.666 Therefore, the error due to discretization also has two parts – one by discretely sampling randomized controls and the other by computing the integrals using finite sums. Szpruch et al., (2024) focuses on the first type of error and Basei et al., (2022) on the second one in their analyses respectively. We address both types of discretization errors in the proof of Theorem 4.1, as well as in Theorems B.3 and B.6 in Appendix B of the full online version of this paper, available at https://arxiv.org/pdf/2407.17226 (which hereafter will be referred to as the “online version”). Therefore, at the $n$th iteration we discretize the interval $[0,T]$ into uniform time intervals of length $\Delta t_{n}$, leading to the following schemes:
@@ 5.3 Analysis of Numerical Results
These experimental results support the theoretical claims and demonstrate the outperformance of our RL-LQ algorithm compared with its model-based counterpart in terms of both the convergence rates of the policy parameters and the regret bounds. [^7] The GitHub repository of the experiments can be accessed at https://github.com/yiliehuang/sublinear-regret-LQ-RL.
@@ A.2 Experiment Setup
  • The projection for $\phi_{1,n}$ is set to be a constant set of $[-2.2,-0.5]$ for computational efficiency. [^8] The projection was originally set to prove the theoretical convergence rate and regret bound. For implementation the theoretical projection grows too slow; instead it could be tuned.
@@
  • The functions $\hat{k}_{1}(t;\boldsymbol{\theta})=1$ and $\hat{k}_{3}(t;\boldsymbol{\theta})=1$ for simplicity, which satisfy the assumptions in Subsection 3.1.999Recall that our results do not dependent on the form of the value function.

### v5

@@ Abstract
footnotetext: Accepted for publication in SIAM Journal on Control and Optimization (2025).
@@ 1 Introduction
Recently, there has been some progress on regret analysis for continuous-time stochastic LQ RL in (Basei et al., , 2022; Szpruch et al., , 2024) that achieves sublinear regrets of their respective algorithms. Both papers assume that the diffusion coefficients are constant (independent of state and control), which is vital for their approaches to work. Again, in essence, these works are model-based because they apply either least-square or Bayesian methods to estimate model parameters and use the corresponding estimation errors to deduce the regret bounds. In particular, there is an intrinsic caveat of this approach applied to LQ problems: when estimating the model parameters, an unidentifiability issue may arise (from optimal control policies being linear feedbacks of the state) and a special exploratory stage has to be carefully designed to avoid it; see e.g., Szpruch et al., (2021); Basei et al., (2022); Szpruch et al., (2024); Xu et al., (2024). In addition, to get one estimate of the model parameters, the required number of episodes and that of timesteps within each episode increase exponentially over the learning process, adding substantial computational and memory costs.For example, given a total $N$ episodes of the learning budget, the algorithm in Basei et al., (2022) suggests to break it into several epoches, with the $k$-th epoch containing $2^{k}$ episodes and $2^{k/2}$ timesteps and to update a policy after having applied it for an epoch of episodes. As a result, although the total regret is proved to be $O(\log N)$, the storage and computational complexity increase exponentially in the number of epoches (which is the number of times of updating the policy parameters).
@@ 2.1 Classical Stochastic LQ Control
We remark that if $x$ is of a higher dimension and/or there is a control running reward, then the optimal policy will depend explicitly on the solution of the differential Riccati equation and hence become time-dependent (Yong and Zhou, , 1999, Chapter 6), for which our current method fails. For example, when $x$ is higher dimensional, the optimal control is $u_{CL}(t,x)=-(\sum_{j=1}^{m}D_{j}^{\top}\frac{\partial V_{CL}(t,x)}{\partial xx}D_{j})^{-1}\left(B^{\top}\frac{\partial V_{CL}(t,x)}{\partial x}+\sum_{j=1}^{m}D_{j}^{\top}\frac{\partial V_{CL}(t,x)}{\partial xx}C_{j}x\right)$, which becomes time-dependent.
@@ 2.2 RL Theory for LQ
where $\bar{\mu}=-(\sum_{j=1}^{m}D_{j}D_{j}^{\top})^{-1}(B+\sum_{j=1}^{m}C_{j}D_{j})$ and $\Sigma=\frac{\gamma}{k_{1}(t)}(\sum_{j=1}^{m}D_{j}D_{j}^{\top})^{-1}$. Detailed derivations of these results follow analogously from those in Wang et al., (2020) for the infinite horizon case. Note that these theoretical results cannot be used to compute the solution of the exploratory problem because the model parameters are unknown, yet they reveal the structure of the solution inherent to LQ RL (i.e. the optimal value function is quadratic in $x$ and optimal stochastic policy is Gaussian) that can be utilized to significantly reduce the complexity of function parameterization/approximation in learning.
@@ 3 A Continuous-Time RL Algorithm
This section presents the steps of designing a continuous-time RL algorithm for solving our LQ problem, including function parameterization, policy evaluation and policy gradient.As will be explained below, policy evaluation is actually not necessary for the specific LQ problem considered in this paper. However, we still include it in the discussion and in the algorithm for future extension to general problems where policy evaluation is generally needed and indeed a crucial step. We will introduce various techniques such as exploration scheduling and projection for deriving the convergence rate of the policy parameter and the regret bound. We will also describe time discretization for final implementation.
@@ 3.1 Function Parameterization
Note that (2.12) suggests that the optimal feedback policy is time-dependent, whose variance depends explicitly on $t$. In our parameterization, the time-dependent variance of the Gaussian policies is replaced by a decaying schedule, called an exploration schedule, of $\phi_{2}$ as a function of the number of iterations, to be presented shortly.An alternative approach is to set a temperature decaying schedule, or equivalently take $\phi_{2}$ as an endogenous, learnable parameter based on the expression (2.12). For example, Szpruch et al., (2024) proposes two different temperature scheduling sequences for their model-based algorithms respectively, which determine the variances of the respective policies in the learning processes that eventually impact the regrets of the algorithms. In this paper, we choose to take a shortcut by taking the entire policy variance as a suitable exogenous tuning sequence (as shown in Theorem 4.1), leaving the temperature $\gamma$ as a constant for simplicity.
@@ 3.5 Discretization
Our approach for continuous-time RL is characterized by carrying out the entire analysis in the continuous-time setting and discretizing time only at the final implementation stage. There are two reasons for discretization. First, note that in the updating rules (3.8) to (3.10), we implicitly assume a continuous sampling of control $u_{n}(t)|x_{n}(t)\sim\mathcal{N}(\phi_{1,n}x_{n}(t),\phi_{2,n})$, which is not practical for actual inplementation. Second, The iterations in (3.8) and (3.9) involve integrals that can be computed only by approximated discretized summations as well as the $\mathrm{d}J$ term that can be approximated by the temporal difference between two consecutive time steps. Therefore, the error due to discretization also has two parts – one by discretely sampling randomized controls and the other by computing the integrals using finite sums. Szpruch et al., (2024) focuses on the first type of error and Basei et al., (2022) on the second one in their analyses respectively. We address both types of discretization errors in the proof of Theorem 4.1, as well as in Theorems B.3 and B.6 in Appendix B of the full online version of this paper, available at https://arxiv.org/pdf/2407.17226 (which hereafter will be referred to as the “online version”). Therefore, at the $n$th iteration we discretize the interval $[0,T]$ into uniform time intervals of length $\Delta t_{n}$, leading to the following schemes:
@@ 5.3 Analysis of Numerical Results
These experimental results support the theoretical claims and demonstrate the outperformance of our RL-LQ algorithm compared with its model-based counterpart in terms of both the convergence rates of the policy parameters and the regret bounds.The GitHub repository of the experiments can be accessed at https://github.com/yiliehuang/sublinear-regret-LQ-RL.
@@ 6 Conclusions
In this paper, the temperature parameter $\gamma$ is fixed and the policy variance $\phi_{2}$ has a prescribed decaying schedule, both independent of data. These two parameters can be regarded as levels of exploration directly controlled by the critic and the actor respectively. A companion paper Huang and Zhou, (2025) considers the same problem but develops different RL algorithms where these two exploration mechanisms are both data driven.
@@ A.2 Experiment Setup
  • The projection for $\phi_{1,n}$ is set to be a constant set of $[-2.2,-0.5]$ for computational efficiency.The projection was originally set to prove the theoretical convergence rate and regret bound. For implementation the theoretical projection grows too slow; instead it could be tuned.
@@
  • The functions $\hat{k}_{1}(t;\boldsymbol{\theta})=1$ and $\hat{k}_{3}(t;\boldsymbol{\theta})=1$ for simplicity, which satisfy the assumptions in Subsection 3.1.Recall that our results do not dependent on the form of the value function.